M is an organic acid with vapour density 75. Percentage composition is C=32% and O=64%. Determine the molecular formula of M acid

Grade:10

1 Answers

Mukhar Jain
7 years ago
Vapour density =75 
therefore Molar mass of acid = 75*2
now it is given it contains 32 %C
64% O
and 4% h
 
we can consider it as in 100 gram of this acid 32 gram is C 64 gram is O and 4 gram is H 
 
now using unitary methode  since 100 gram contain 32 gram C ------>1 gram contain 32/100 gram C
and 150 gram conatin 150*(32/100)=48gram 
there fore 4 moles of carbon 
 
simillarly 150*64/100=96 ====6 moles O and 6 gram h 
 
now the acis id C4 O6 H6